how can you graph y=2x-8 nd y=2/3x on a graph, using y=mx+b form?

Answer by kkasko(46)

how can you graph y=2x-8 nd y=2/3x on a graph, using y=mx+b form?
The mx is the slope which means it designs what the line will look like for instance in y=2x-8, the 2x is the slope and it has a rise of 2 and a run of 1. The rise is on the y axis and the run is on the x- axis. So the slope is up 2 right 1. The b is the y intercept and is the point (0,-8) which is on the y axis. In order to graph this equation draw a point on (0,-8) and go up 2 right 1. And do the opposite. Down 2 left one once you have drawn as far as possible.
To plot the next equation the y intercept is (0,0). Plot that on the y axis and it has a rise of 2 and a run of 3 so go up 2 and right 3 till you have drawn as far as possible and then do the opposite. Go down 2 and left 3 from the point (0,0) that you haven't drawn on yet.
The points should have intersected on (6,4)
X=6
Y=4